A Toblerone bar comes in packaging that forms a triangular prism with equilateral triangles at both ends. If the volume of the package is 20cm3, determine the dimensions of the package that give a minimum surface area. V = 1/2bhl, S = bh + 3bl 


1
Expert's answer
2021-03-26T08:46:52-0400

Given-

A Toblerone bar comes in packaging that forms a triangular prism with equilateral triangles at both ends.

So, its length and breadth is equal means "b=l"

"V=20cm^3" "V=\dfrac{1}{2bhl}"

A.T.Q "S=bh+3bl"

"\Rightarrow" "S=\dfrac{1}{2Vl}+3l^2" "(" "V=\dfrac{1}{2bhl}," "b=l)"

"\Rightarrow" differentiate the above equation with respect to "l"

"\Rightarrow" "\dfrac{dS}{dl}=\dfrac{-l^2}{2V}+6l^3"

put this equation equal to zero for minimum surface area

"\Rightarrow" "l=\dfrac{1}{12V} =0.0046 cm"

which is equal to "b=0.0046 cm"

then "h=1181.47cm" "(h=\dfrac{1}{2Vbl})"